OpenStudy (anonymous):

20 x 3/5 = ? Simplest form

OpenStudy (anonymous):

its 12. divide 3 by 5 and then multiply that by 20.

OpenStudy (some.random.cool.kid):

yep

OpenStudy (some.random.cool.kid):

exactly

OpenStudy (anonymous):

1/5 x 8 = ? simplest form

OpenStudy (anonymous):

1.6

OpenStudy (anonymous):

it is 1.6. Divide 1 and 5 then multiply that by 8

OpenStudy (skullpatrol):

8 can be written as 8/1 and then multiply the numerators and multiply the denominators.

OpenStudy (skullpatrol):

$$\Huge\frac{a}{b}\cdot\frac{c}{d}=\frac{ac}{bd}$$